# The Illusion of Progress? A Critical Look at Test-Time Adaptation for Vision-Language Models

### v1

@@ Abstract
Test-time adaptation (TTA) methods have gained significant attention for enhancing the performance of vision-language models (VLMs) such as CLIP during inference, without requiring additional labeled data. However, current TTA researches generally suffer from major limitations such as duplication of baseline results, limited evaluation metrics, inconsistent experimental settings, and insufficient analysis. These problems hinder fair comparisons between TTA methods and obscure their practical strengths and weaknesses. To address these challenges, we introduce TTA-VLM, a comprehensive benchmark for evaluating TTA methods on VLMs. Our benchmark implements 8 episodic TTA and 7 online TTA methods within a unified and reproducible framework, and evaluates them across 15 widely used datasets. Unlike prior studies focused solely on CLIP, we extend the evaluation to SigLIP—a model trained with a Sigmoid loss—and include training-time tuning methods such as CoOp, MaPLe, and TeCoA to assess generality. Beyond classification accuracy, TTA-VLM incorporates various evaluation metrics, including robustness, calibration, out-of-distribution detection, and stability, enabling a more holistic assessment of TTA methods. Through extensive experiments, we find that 1) existing TTA methods produce limited gains compared to the previous pioneering work; 2) current TTA methods exhibit poor collaboration with training-time fine-tuning methods; 3) accuracy gains frequently come at the cost of reduced model trustworthiness. We release TTA-VLM to provide fair comparison and comprehensive evaluation of TTA methods for VLMs, and we hope it encourages the community to develop more reliable and generalizable TTA strategies. The code is available in https://github.com/TomSheng21/tta-vlm.
@@ 1 Introduction
Vision-language models (VLMs) [1, 2, 3, 4, 5], such as CLIP [1] learn to align visual and textual representations in a shared embedding space, enabling effective performance on a wide range of multi-modal tasks. These models achieve remarkable performance in tasks such as image classification [1, 6, 7], multi-modal retrieval [8, 9, 10], and semantic segmentation [11, 12, 13, 14]. Especially, the introduction of natural language makes VLMs have better flexibility and generalization than traditional visual classification models. To further enhance the performance of VLMs, numerous training-time approaches [6, 15, 16, 17, 18] have been proposed, introducing effective optimization solutions and inference paradigms.
In contrast to training-time approaches, test-time adaptation (TTA) methods [19, 20, 21, 22] have attracted increasing attention for their ability to enhance the performance of VLMs during inference without requiring additional annotated data. Depending on the type of test data, TTA methods can be broadly categorized into episodic [21, 23, 24] and online [25, 26, 27, 28] TTA strategies. Episodic TTA adapts the model to a single test sample, exploring the information inside the sample to improve the model’s prediction. Online TTA aims to process a stream of test data, utilizing historical knowledge in the previously seen data to update model behavior.
@@ 2.2 Test-Time Adaptation Methods
Our benchmark contains 8 episodic TTA methods and 7 online TTA methods. We provide a brief introduction here, and a detailed description can be found in the supplementary material.
@@ 5 Trustworthiness and Stability from Accuracy-oriented Adaptation
CLIP [1] 0.00 0.16 0.00 0.00 0.00 0.00 18.06 0.00 0.00 0.00 0.00 1.82
TPT [21] 0.00 0.24 0.00 0.00 0.00 0.00 18.93 0.00 0.06 0.00 0.00 1.92
C-TPT [23] 0.00 0.16 0.00 0.00 0.00 0.00 18.12 0.00 0.00 0.00 0.00 1.83
RLCF [38] 0.00 0.28 0.00 0.00 0.00 0.00 17.46 0.00 0.30 0.00 0.00 1.80
MTA [39] 25.15 67.30 46.31 10.25 20.34 27.46 20.19 25.28 13.59 0.32 29.50 26.05
ZERO [40] 38.93 76.71 58.16 27.80 36.50 41.09 20.07 42.49 27.90 4.91 41.13 37.68
TTL [41] 4.26 22.43 9.46 11.96 11.94 9.74 18.54 3.95 5.61 0.01 4.12 9.78
TPS [42] 0.00 0.20 0.00 0.00 0.00 0.00 19.71 0.00 0.00 0.00 0.00 1.99
R-TPT [24] 38.61 76.11 56.75 27.57 37.27 41.52 19.53 42.65 27.90 5.30 41.13 37.57
@@ Appendix A Introduction of TTA Methods
$\min_{\theta_{\text{p}}}\ \mathcal{L}=\mathcal{L}_{TPT}+\mathcal{L}_{CTPT}=\mathcal{L}_{TPT}+\sum_{k=1}^{C}{||\bar{t}-t_{k}||},$ (2)
@@ Appendix B Introduction of Datasets
We introduce the detailed information of the datasets included in the benchmark in Table 12.

### v2

@@ Abstract
Test-time adaptation (TTA) methods have gained significant attention for enhancing the performance of vision-language models (VLMs) such as CLIP during inference, without requiring additional labeled data. However, current TTA researches generally suffer from major limitations such as duplication of baseline results, limited evaluation metrics, inconsistent experimental settings, and insufficient analysis. These problems hinder fair comparisons between TTA methods and make it difficult to assess their practical strengths and weaknesses. To address these challenges, we introduce TTA-VLM, a comprehensive benchmark for evaluating TTA methods on VLMs. Our benchmark implements 8 episodic TTA and 7 online TTA methods within a unified and reproducible framework, and evaluates them across 15 widely used datasets. Unlike prior studies focused solely on CLIP, we extend the evaluation to SigLIP—a model trained with a Sigmoid loss—and include training-time tuning methods such as CoOp, MaPLe, and TeCoA to assess generality. Beyond classification accuracy, TTA-VLM incorporates various evaluation metrics, including robustness, calibration, out-of-distribution detection, and stability, enabling a more holistic assessment of TTA methods. Through extensive experiments, we find that 1) existing TTA methods produce limited gains compared to the previous pioneering work; 2) current TTA methods exhibit poor collaboration with training-time fine-tuning methods; 3) accuracy gains frequently come at the cost of reduced model trustworthiness. We release TTA-VLM to provide fair comparison and comprehensive evaluation of TTA methods for VLMs, and we hope it encourages the community to develop more reliable and generalizable TTA strategies. The code is available in https://github.com/TomSheng21/tta-vlm.
@@ 1 Introduction
Vision-language models (VLMs) [1, 2, 3, 4, 5], such as CLIP [1], learn to align visual and textual representations in a shared embedding space, enabling effective performance on a wide range of multi-modal tasks. These models achieve remarkable performance in tasks such as image classification [1, 6, 7], multi-modal retrieval [8, 9, 10], and semantic segmentation [11, 12, 13, 14]. Especially, the introduction of natural language makes VLMs have better flexibility and generalization than traditional visual classification models. To further enhance the performance of VLMs, numerous training-time approaches [6, 15, 16, 17, 18] have been proposed, introducing effective optimization solutions and inference paradigms.
In contrast to training-time approaches, test-time adaptation (TTA) methods [19, 20, 21, 22, 23] have attracted increasing attention for their ability to enhance the performance of VLMs during inference without requiring additional annotated data. Depending on the type of test data, TTA methods can be broadly categorized into episodic [21, 24, 25] and online [26, 27, 28, 29] TTA strategies. Episodic TTA adapts the model to a single test sample, exploring the information inside the sample to improve the model’s prediction. Online TTA aims to process a stream of test data, utilizing historical knowledge in the previously seen data to update model behavior.
@@ 2.2 Test-Time Adaptation Methods
Our benchmark contains 8 episodic TTA methods and 7 online TTA methods. We provide a brief introduction here; a detailed description and the introduction of TTA methods for other tasks (e.g., segmentation) can be found in the supplementary material.
@@ 4 Collaboration with Training-time Fine-tuning Methods
Remark. Regarding whether the deployer needs to introduce a training time phase before the test-time method, we have the following suggestions: When the labeled training data and test data share the same or highly overlapping categories, training-time methods can lead to significant gains. For example, fine-tuning on labeled ImageNet data yields noticeable improvements on downstream tasks such as ImageNet-V2 (4-shot ImageNet tuning w/o TTA: 51.5%->55.6% on ImageNet-V2). Also, when robustness or other specific properties are required, a supervised train-time stage (e.g., adversarial fine-tuning) can provide a more robust initialization for TTA. Conversely, if the labeled data is not strongly correlated with the test distribution, supervised fine-tuning tends to generalize poorly (4-shot ImageNet tuning w/o TTA: 55.8%->56.2% on fine-grained dataset). In such scenarios, we recommend applying TTA directly, as it can better adapt to downstream tasks.
@@ 5 Trustworthiness and Stability from Accuracy-oriented Adaptation
CLIP [1] 0.00 0.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.02
TPT [21] 0.00 0.24 0.00 0.00 0.00 0.00 0.00 0.00 0.06 0.00 0.00 0.03
C-TPT [24] 0.00 0.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.02
RLCF [39] 0.00 0.28 0.00 0.00 0.00 0.00 0.00 0.00 0.30 0.00 0.00 0.06
MTA [40] 25.15 67.30 46.31 10.25 20.34 27.46 2.04 25.28 13.59 0.32 29.50 24.24
ZERO [41] 38.93 76.71 58.16 27.80 36.50 41.09 8.88 42.49 27.90 4.91 41.13 36.56
TTL [42] 4.26 22.43 9.46 11.96 11.94 9.74 3.42 3.95 5.61 0.01 4.12 8.26
TPS [43] 0.00 0.20 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.02
R-TPT [25] 38.61 76.11 56.75 27.57 37.27 41.52 9.00 42.65 27.90 5.30 41.13 36.52
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
Funding in direct support of this work was provided by the National Natural Science Foundation of China under Grants (62276256, U2441251), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Young Scientists Fund of the State Key Laboratory of Multimodal Artificial Intelligence Systems (ES2P100117).
The authors declare that they have no competing interests.
@@ Appendix A Introduction of TTA Methods
[…]
@@
TTA methods beyond classification. Although our benchmark focuses on classification tasks, several recent works explore TTA in broader visual domains such as segmentation, detection, anomaly localization, and video understanding. We briefly introduce their core ideas below.
Among episodic TTA approaches, AnoCLIP [70] proposes a zero-shot anomaly localization framework that enhances CLIP’s patch-level alignment and introduces a lightweight test-time adapter for refining localization. T3AL [71] performs episodic test-time adaptation for zero-shot temporal action localization by generating video-level pseudo-labels and refining temporal regions using a self-supervised strategy. DTS-TPT [72] extends episodic test-time prompt tuning to video activity recognition, aligning text prompts with multi-scale temporal features for improved zero-shot accuracy. CLIP-DIY [73] achieves training-free open-vocabulary semantic segmentation by aggregating CLIP’s patch-level predictions with unsupervised localization guidance at test time. CLIPtrase [74] recalibrates CLIP’s patch correlations to improve local feature discrimination, boosting training-free semantic segmentation performance. VocAda [75] introduces a plug-and-play vocabulary adapter for open-vocabulary detection that refines user-defined labels through image captioning and noun selection at test time. As for online TTA, TTCS [76] adapts the Segment Anything Model for medical image segmentation using CLIP-guided prompt generation and adaptive self-training. TEST-V [77] combines prompt-based and support-set adaptation for zero-shot video classification, tuning support samples dynamically to enhance temporal consistency. These methods show that TTA is expanding beyond classification toward dense and video tasks, highlighting the versatility of adaptation-based paradigms for broader vision-language applications.
@@
(a) Caltech101
(b) Pets
(c) Cars
(d) Flowers
(e) Food101
(f) Aircraft
(g) SUN397
(h) DTD
(i) EuroSAT
(j) UCF101
(k) ImageNet
(l) ImageNet-A
(m) ImageNet-V2
(n) ImageNet-R
(o) ImageNet-S
Figure 3: Examples from the 15 datasets used in TTA-VLM, covering objects, animals, scenes, actions, textures, and cross-domain visual tasks.
@@ Appendix B Introduction of Datasets
We introduce the detailed information of the datasets included in the benchmark in Table 12 and provide representative examples in Figure 3.
@@ Appendix D Detailed Experimental Results
Visualization of visual features. We employ t-SNE to visualize the visual representations learned by CLIP (ViT-B/32) [1] and TTL [42] on a subset of the UCF101 dataset. As shown in Figure 4, the features generated by TTL exhibit more compact and well-separated clusters compared to the original zero-shot features from CLIP, with fewer ambiguous samples lying near decision boundaries. This indicates that TTL produces more discriminative and robust representations, consistent with its superior classification accuracy and adversarial robustness observed in quantitative experiments.
(a) (1) CLIP (ViT-B/32)
(b) (2) TTL
Figure 4: t-SNE visualization of visual features from CLIP and TTL on a subset of the UCF101 dataset.